OpenStudy (anonymous):

A cheetah's sprinting speed is 61 miles per hour less than three times a zebra's running speed. A cheetah can sprint 68 miles per hour. Find the zebra's running speed.

OpenStudy (anonymous):

OK... This problem talks about a cheetah's speed and a zebra's speed. Does the problem say a cheetah's speed IS or a zebra's speed IS

OpenStudy (anonymous):

Only that a cheetah can sprint 68 miles per hour

OpenStudy (anonymous):

If the zebra can sprint at a speed of x, then the cheetah can sprint at a speed of 3x-61. If the cheetah sprints at a speed of 68mph, then 68=3x-61.

OpenStudy (anonymous):

To solve for the zebra's speed, you solve for x.

OpenStudy (anonymous):

Read the first 5 or 6 words of the problem. Stay with me here, I will teach you how to do these... It says a cheetah's sprinting speed IS.... do you see that

OpenStudy (anonymous):

you can't solve for x because by then the answer will be 43 which is < 61, they said that the zebra is 3 times faster :)

OpenStudy (anonymous):

3z - 61 is the formula, if cheeta = z = 61 then zebra = 3(61)

OpenStudy (anonymous):

if cheetah = 68 mi/hr then: zebra = z = 3(68) = 204 mi/hr

OpenStudy (anonymous):

blexting, you may find purchase in starting a thread to explain the process and having questions asked there if that was your intention. I'll quit passing out answers.

OpenStudy (anonymous):

Correct me if I'm wrong ^_^

OpenStudy (anonymous):

The cheetah sprints at 61mph less than three times the zebras speed. That shows that the cheetah is less than three times as fast as the zebra, not that the zebra is three times as fast as a cheetah.

OpenStudy (anonymous):

no the cheetah is 3 times faster .

OpenStudy (anonymous):

Lil B are you still there?

OpenStudy (anonymous):

explain blexting and he'll come here and check later ^_^

OpenStudy (anonymous):

Right. Which means for a zebra of speed Z, the cheetah is a speed of 3Z. However, the cheetah is 61mph less than 3Z, so it runs at a speed of 3z-1.

OpenStudy (anonymous):

dear, the zebra will run at 3z, while the cheetah will have a speed that is less than that of a zebra

OpenStudy (anonymous):

But that is where you are wrong. The cheetah's sprinting speed is some value less than three times that of the zebra, not that the zebra is in any way three times as fast as a cheetah.

OpenStudy (anonymous):

After reading what everyone as written, Im still a bit hazey on how to express my equation.

OpenStudy (anonymous):

lol, don't worry Lil B , we're just discussing ^_^

OpenStudy (anonymous):

OK.. I just noticed that you had problems on one like this before so I thought I would teach you how to do these.

OpenStudy (anonymous):

I do greatly appreciate everyone's help!

OpenStudy (anonymous):

blexting, please explain it to him if you can :)

OpenStudy (anonymous):

A cheetah's sprinting speed is 61 miles per hour less than three times a zebra's running speed. This statement converts to an algebraic expression: C=3Z-61. The 3Z represents "three times that zebra's speed" A cheetah can sprint 68 miles per hour. Find the zebra's running speed. So we know that C=61. Then our formula becomes: 61=3Z-61. And now it's an algebra problem.

OpenStudy (anonymous):

C=68, but yes.

OpenStudy (anonymous):

Ack. Sorry about my typos. Doesn't happen to me when I actually write the problem out. :)

OpenStudy (anonymous):

Since C=68, then the formula we discovered C=3Z-61 becomes 68=3Z-61

OpenStudy (anonymous):

Lil B, Blexting is explaining the process that we use to get these formulas in another thread that I started.

OpenStudy (anonymous):

Ok! thanks guys! @jcntrl: Im with you!